OpenStudy (anonymous):

What is the maximum volume of a pyramid that can fit inside a cube with sides that are 18 cm long?

OpenStudy (anonymous):

cube volume = a*a*a=a^3 pyramid=L*W*H*1/3 L=a, W=a, H=a so pyramid=a^3/3=(18)^3/3=1944 cm^3